[Brahms-l] Shift report 20050428 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Apr 28 2005 - 07:54:28 EDT
Shifter leader was sk
Shifters were 


1. 
No beam till 5am. Machine has cryo lead flow problem.
2. 
FS 8 deg D1 244A, MSR 45deg D5 1050A
Run   Tr2   Tr3
14319 0     177302
14322 0     174249
Pedestal Runs:
14317 => CAMAC
14318 +> TPC
3.
Problem with magnet D1 fixed.
No HV for:
TOF-II :
MCh0,  MCh8,  MCh17,  MCh18,   MCh19
H2:
No connection to mainframe.
TOFW:
During shifts problem with connection, after restart BRAVO seems to be o.k.

Current on T5 A2y seems to be high:
A2x 0.38 A2y 1.41 A2u 0.44 A2v 0,49


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Apr 28 07:55:06 2005

This archive was generated by hypermail 2.1.8 : Thu Apr 28 2005 - 07:55:25 EDT